OpenStudy (anonymous):

P = M - N over 2. Solve for N Please and Thank yous <3

OpenStudy (anonymous):

n = m-2p

hero (hero):

Juanita, how come you dont' post your steps?

OpenStudy (anonymous):

I need los steps </3

OpenStudy (anonymous):

P = (M - N)/2 2P = M - N 2P - M = -N -(2P - M) = N -2P + M = N

hero (hero):

los..wow I haven't seen that in a while

hero (hero):

Does los = those?

hero (hero):

or does it just mean a plural form of the?

OpenStudy (anonymous):

The plural form of the